Larry, a 40 year old salesperson, earns $90,000 per year and plans to work until age 65. He is married to Joan and has 2 children. He expects his annual salary increases to be 3%, and the inflation rate to be 3%, and their average tax bracket (state and federal) is 25%. Larry estimates that 15% of his after-tax income is used for personal consumption. Based on the Human Life Value approach to life insurance needs analysis, how much life insurance should Larry purchase?
Answer #1

Please do Upvote if you are served. Feel free to reach out in the comments

Cheers!!!

Answer:

Given Larry earns $90,000

After tax earnings = 90000*(1-0.25) = $67500

After personal consumption = 675000 (1-0.15) = $57,375

Let us calculate the future value with these conditions and later plough back the FV to PV to consider how much insurance to buy

Now payments, PMT = $57,375, n = 25 years, i =3%(salary increase)

=> Calculating FV for this payments = FV(0.03,25,-57375) (Using excel) = $2,091,850.29

Now currently, the insurance premium can be found by calculating the PV for $2,091,850.29 by having 3% inflation rate

=> PV(0.03,25,,2091850.29) = $999079.35